OpenStudy (abbycross167):

Can someone please help me answer this question? I have an idea of what it could be but I'm not sure... Question~ graph the following inequalities. The accuracy of the line, if the line is dashed or solid, and which side you shade are important components. Make your points in yellow, the line in blue and shade in red. x + y > -4 NOTE* The > has a line under it

OpenStudy (abbycross167):

So would X= -4+ -1y?

OpenStudy (abbycross167):

@phi could you try to help me please?

OpenStudy (studygurl14):

move the x to the other side to get y > -x - 4 Now graph that.

OpenStudy (phi):

**The > has a line under it*** if you can't type in \( \ge\) you can use >= it means greater than or equal

OpenStudy (phi):

if you have a >= or <= you will make the line solid if it is just > or < then use a dashed line your problem will use a solid line

OpenStudy (abbycross167):

yes sir/ ma'am, I know that part

OpenStudy (phi):

to plot the line, the normal slope intercept form is y= mx + b in other words, write your equation x+y >= -4 by adding -x to both sides and simplifying. you get y >= -x - 4 now put in x=0. what do you get for y ?

OpenStudy (abbycross167):

y=-4?

OpenStudy (phi):

yes, so plot (0, -4) do you know how?

OpenStudy (abbycross167):

Hmm I'm a little confused, would you mind showing me the points I should plot? I can draw the line and everything else, but I don't know where the points should be

OpenStudy (phi):

First, let me fix the 2nd point I said pick x=1 and figure out y y= -x-4 so y= -1-4= -5 and the point is (1, -5)

OpenStudy (phi):

to plot (0,-4) find the origin (where x and y cross) now count straight down 4 steps and put a yellow dot there. can you do that ?

OpenStudy (abbycross167):

 Like that? @phi I know the dot isn't yellow I will fix that in my lesson

OpenStudy (phi):

yes looks good now for (1, -5) go to the origin and count to the right 1 step, then down 5 steps

OpenStudy (phi):

you should be 1 to the right and 1 below the first point

OpenStudy (abbycross167):